OpenStudy (anonymous):

Yasmin is training for a marathon, and each month she increases her running distance. During her 4 months of training, her distances were 3, 6, 12 , and 24 miles. What is the rule the sequence follows? Add 3 miles to the previous number. Divide the previous number by 2. Multiply the previous number by 2. Subtract 3 miles from the previous number.

OpenStudy (anonymous):

the rule is multiplication of the previous number by 2: 3 * 2 = 6 6 * 2 = 12 12 * 2 = 24